### v1

# Spontaneous emission of a three-level artificial atom in a one-dimensional open waveguide

Abstract
We study the dynamical and spectral characteristics of a quantum three-level ladder system, interacting with a continuous electromagnetic field in one-dimensional open waveguide. Common realization of such systems is a waveguide QED setup  — a superconducting artificial atom (transmon), coupled to an open microwave transmission line. We derive an analytical solution for spontaneous emission of initially excited atom, and use it to study the probability of state detection and spectral density of output photon states. We find that for strong coupling of transmon to a waveguide emitted photons show correlation in frequency and can have the same energies, even if the three-level system is anharmonic.
waveguide QED, photon-photon correlation, artificial atom
pacs:
84.40.Az,  84.40.Dc,  85.25.Hv,  42.50.Dv, 42.50.Pq
I Introduction
Waveguide quantum electrodynamics is a modern and rapidly developing field of research that studies the interaction of atoms with photons in confined geometries of waveguides Roy17 ; Gu17 ; Sheremet23 ; Astafiev10 . It has a vast spectrum of applications in quantum optics and photonics, as well as in quantum technologies such as quantum communication and quantum information processing. In contrast with cavity QED Raimond01 ; Haroche06 ; Kibis11 ; Savenko12 ; Kibis13 and circuit QED Blais04 ; Wallraff04 ; Blais21 that are focused more on resonant interaction of atoms and nanostructures with single-mode fields in cavities or resonators, waveguide QED describes interaction with a continuum of photon modes in open waveguides.
Three-level systems play an important role in quantum optics and photonics, and have recently received broad attention in waveguide QED as well. Interaction of three-level emitters (3LE) with photon fields can induce some interesting and important effects, such as electromagnetically induced transparency (EIT) Abdumalikov10 , Autler-Townes splitting Sillanpaa09 ; Anisimov11 , two-photon resonance fluorescence Gasparinetti19 ; Ngaha25 ; Gunin25 , giant cross-Kerr effect Hoi13a , photon-state purification Lopes24 , and many others. Since an excited three-level system can emit two photons via spontaneous decay or scattering interaction, it can be potentially used to induce photon-photon correlations. Creation of correlated photon pairs due to spontaneous decay or scattering on a single atom in 1D waveguides was studied in Shen07a ; Shen07b ; Roy11 ; Zheng12 ; Chen12 ; Li15 ; Sathyamoorthy16 ; Gasparinetti17 ; Gao24 .
A typical realization of a 3LE interacting with a continuum of photons in 1D waveguides is superconducting artificial atoms Kjaergaard20 ; Krantz19 , directly or side-coupled to a microwave transmission line Astafiev10 ; Blais21 . The most common type of a three-level system in this playground is a transmon qubit Koch07 . A transmon has a ladder-type structure of levels, meaning that transitions from a lower to a higher level can only happen via a middle level. Although transmons are usually used for quantum information applications as regular qubits, i.e., two-level systems, the third level can have an important practical role in waveguide QED setups Hoi11 ; Lu21 ; Sultanov25 , and, of course, has wide quantum optics applications (see Hoi13b and references from the previous paragraph).
In this work we study the spontaneous emission of an initially excited three-level system, coupled to a 1D open waveguide. In contrast with more traditional tasks like interaction of radiation with single qubits Shen05 ; Chen11 ; Wang11 or qubit chains Fang14 ; Greenberg15 ; Liao15 ; Greenberg21 ; Greenberg23 , that are usually performed in the single-photon limit, here we employ the two-excitation subspace. Our approach is based on the real-space description of photon fields. We derive an analytical solution of the non-stationary Schrödinger equation and find the probability amplitudes for the complete wavefunction of the system. Using this solution, we analyze probabilities to detect certain states of the system, and the spectral density of output two-photon fields. For a weak radiative decay rate we find that each of the two photons has a distinct frequency, corresponding to the two transitions of the 3LE, which is in good agreement with the experimental work Gasparinetti17 . However, if the radiative decay is approximately ten times higher (which is achievable in current setups), we find that frequency correlation appears, which can result in the creation of two identical photons with equal energies. This correlation is more pronounced for low anharmonicity of the 3LE (i.e., when the levels are almost equidistant). We show that depending on the ratio of decay rates from the upper and middle levels, the spectral density for identical photons can have two maxima: on resonance with the lower two levels, or resonant with a smaller frequency which differs from the first frequency by half the anharmonicity parameter.
The paper is organized as follows. In Section II we give a general description of the system under study and derive the main equations for the wavefunction’s amplitudes. In Section III we derive a general analytical solution in real-space. In Section IV we examine the case of a fully-excited three-level atom, both in real-space and in frequency representations. The main results are placed in this section, including analysis of state-detection probabilities, two-photon emission spectra and creation of correlated photon pairs that share the same frequencies. The conclusion of our work, as well as future prospects, are reflected in Section V. Some mathematical calculations are included in the Appendix.
II Real-space description of the transmon-waveguide system
We consider a system of a ladder-type three-level emitter (3LE), located at the point $x=0$ in a one-dimensional waveguide (see Fig. 1a). In the real-space representation Shen07b ; Shen09 and under rotating wave approximation it can be described by the following Hamiltonian (here and later in the text we assume that $\hbar=1$):
$\begin{gathered}H=\Omega_{1}\left|e\right\rangle\left\langle e\right|+\Omega_{2}\left|f\right\rangle\left\langle f\right|-i\upsilon_{g}\int\limits_{-\infty}^{+\infty}dx\;a^{\dagger}(x)\frac{\partial}{\partial x}a(x)\\ +\int\limits_{-\infty}^{+\infty}dx\;V_{1}\delta(x)\left({a^{\dagger}(x)\left|g\right\rangle\left\langle e\right|+a(x)\left|e\right\rangle\left\langle g\right|}\right)\\ +\int\limits_{-\infty}^{+\infty}dx\;V_{2}\delta(x)\left({a^{\dagger}(x)\left|e\right\rangle\left\langle f\right|+a(x)\left|f\right\rangle\left\langle e\right|}\right).\end{gathered}$ (1)
Here states $\left|g\right\rangle,\left|e\right\rangle$ and $\left|f\right\rangle$ are ground, first-excited and second-excited states of the transmon. Corresponding resonant frequencies of first and second excited states are $\Omega_{1},\Omega_{2}$, and $V_{1},V_{2}$ are their coupling constants. Energy levels and corresponding frequencies are shown in Fig. 1b. $\upsilon_{g}$ is a group velocity of photons, $\delta(x)$ is a Dirac delta-function, and $a^{\dagger}(x)$ ($a(x)$) is a creation (annihilation) operator in x-representation, which creates (destroys) photon in $x$ position of the waveguide.
Throughout of this paper we will refer to 3LE as transmon or (artificial) atom, but keep in mind that our approach is universal and can be related to other physical realizations of three-level ladder-type systems.
Figure 1: (a) System under study - three-level emitter (3LE), coupled to the center of one-dimensional waveguide at point $x=0$. Coupling of first and second excited states are $V_{1}$ and $V_{2}$, respectfully. Excited atom can emit photons through spontaneous decay, which are propagate only forward due to usage of chiral approximation (see text for details). (b) Energy levels diagram of the bare transmon and corresponding resonant frequencies.
Hamiltonian (1) is written in the so-called continuous limit Blow90 , meaning that instead of discrete summation over $x$ (or over modes $k$), we have continuous integration. In this limit, commutation relation between photon operators is as follows:
$\left[a(x),a^{\dagger}(x^{\prime})\right]=\delta(x-x^{\prime}),$ (2)
hence the photon operators, usually dimensionless for discrete modes, in continuous limit have dimension of inverse square root of coordinate, $a(x)\sim 1/\sqrt{x}$. The dimension of coupling constants $V_{1},V_{2}$, correspondingly, is $V\sim\omega\sqrt{x}$.
The first two terms in (1) correspond to a free-atom and free-field, and the following two terms describe the interaction of photon fields with first and second excited states of transmon, respectively. Note that there is no direct transition between ground and second-excited state $\left|f\right\rangle\leftrightarrow\left|g\right\rangle$. Hamiltonian (1) is practically Janes-Cumming Hamiltonian, extended for three-level system. It conserves number of excitations in the system. In this work we adhere two excitations due to two possible transitions of the transmon.
Since our atom have three levels, wave function should be written in a two-excitation subspace as follows:
$\begin{gathered}\left|\Psi(t)\right\rangle=\alpha(t)e^{-i\Omega_{2}t}\left|f,0\right\rangle+\int\limits_{-\infty}^{+\infty}dxe^{-i\Omega_{1}t}\beta(x,t)a^{\dagger}(x)\left|e,0\right\rangle\\ +\frac{1}{\sqrt{2}}\iint\limits_{-\infty}^{+\infty}dx_{1}dx_{2}\;\gamma(x_{1},x_{2},t)a^{\dagger}(x_{1})a^{\dagger}(x_{2})\left|g,0\right\rangle,\end{gathered}$ (3)
where first probability amplitude $\alpha(t)$ corresponds to fully excited transmon in state $\left|f\right\rangle$ and no photons in a waveguide, second amplitude $\beta(x,t)$ corresponds to transmon excited to $\left|e\right\rangle$ state and one photon at position $x$, and final amplitude $\gamma(x_{1},x_{2},t)$ contains non-excited atom and two photons in $x_{1}$ and $x_{2}$ positions. State $\left|0\right\rangle$ designates vacuum Fock state. Coefficient $1/\sqrt{2}$ in last term is a normalization constant for two-photon state in Fock basis.
Because of continuous limit Blow90 , probability amplitudes with photons have dimension of inverse square root of coordinate: $\beta(x,t)\sim 1/\sqrt{x}$, and $\gamma(x_{1},x_{2},t)\sim 1/x$ due to presence of two photons.
We should also note that wavefunction (3), as well as Hamiltonian (1), describe photon propagation only in forward direction. This is so called chiral approximation, and it can be performed by separating Hamiltonian for odd and even modes Shen07b . Chiral, or non-reciprocal, waveguide can be realized experimentally by creating suppression of photon propagation in backward direction Lodahl17 ; Guimond20 . We are mainly interested in the case of spontaneous decay of initially excited 3LE, which should be symmetrical to 3LE placement. So, in order to simplify analytical derivation, we choose to consider only forward propagation.
Using standard non-stationary Schrödinger equation $i\partial_{t}\left|\Psi\right\rangle=H\left|\Psi\right\rangle$ with Hamiltonian (1) and wave function (3), we derive the equations for all probability amplitudes:
$\frac{d}{dt}\alpha(t)=-iV_{2}{e^{i(\Omega_{2}-\Omega_{1})t}}\beta(0,t),$ (4)
$\begin{gathered}\frac{d}{dt}\beta(x,t)+\upsilon_{g}\frac{\partial}{\partial x}\beta(x,t)=-iV_{2}\delta(x){e^{i(\Omega_{1}-\Omega_{2})t}}\alpha(t)\\ -i\frac{2}{\sqrt{2}}V_{1}{e^{i\Omega_{1}t}}\gamma(x,0,t),\end{gathered}$ (5)
$\begin{gathered}\frac{d}{dt}\gamma(x_{1},x_{2},t)+\upsilon_{g}\frac{\partial}{\partial x_{1}}\gamma(x_{1},x_{2},t)+\upsilon_{g}\frac{\partial}{\partial x_{2}}\gamma(x_{1},x_{2},t)\\ =-\frac{i}{\sqrt{2}}V_{1}{e^{-i\Omega_{1}t}}\left[{\delta(x_{1})\beta(x_{2},t)+\delta(x_{2})\beta(x_{1},t)}\right].\end{gathered}$ (6)
In the last two equations we use the fact that photons are fundamentally indistinguishable, which in our case can be stated as $\gamma(x_{1},x_{2},t)=\gamma(x_{2},x_{1},t)$. All further calculations are based on equations (4-6).
III Probability amplitudes in real-space
To find explicit expressions for all three amplitudes of wave function, we use method, incorporated in Mukhopadhyay24 for calculation of quantum Rabi oscillation dynamics and also employed in Lopes24 for photon state-purification via 3LE interaction. We have already used this approach in Chuikin24 for description of stimulated emission in waveguide QED, which is very similar task to that we solve here.
First, let’s define a Fourier transform for amplitudes with photonic part, which allow us to switch from real-space to k-space of frequency modes:
$\beta(k,t)$ $=\frac{1}{\sqrt{2\pi}}\int\limits_{-\infty}^{+\infty}{dx\;}{e^{-ikx}}\beta(x,t),$ (7a)
$\gamma(k_{1},k_{2},t)$ $=\frac{1}{2\pi}\iint\limits_{-\infty}^{+\infty}{dx_{1}dx_{2}\;}{e^{-ik_{1}x_{1}}}{e^{ik_{2}x_{2}}}\gamma(x_{1},x_{2},t).$ (7b)
Inverse transformation is done by changing sign of exponents.
Here and further we assume that dispersion is linear, $k=\omega/\upsilon_{g}$, which is well justified if the atom and photon resonant frequencies are far from cut-off frequency of the waveguide Shen07b ; Shen09 .
By using (7) in (6) together with integral representation of delta function as $\delta(x)=1/2\pi\int dk\;e^{ikx}$, we obtain equation for two-photon amplitude in k-space:
$\begin{gathered}\frac{d}{dt}\gamma(k_{1},k_{2},t)+i\upsilon_{g}(k_{1}+k_{2})\gamma(k_{1},k_{2},t)\\ =-\frac{i}{\sqrt{2}\sqrt{2\pi}}V_{1}{e^{-i\Omega_{1}t}}\left[{\beta(k_{2},t)+\beta(k_{1},t)}\right].\end{gathered}$ (8)
Formal solution of this equation is:
$\begin{gathered}\gamma(k_{1},k_{2},t)=-\frac{i}{\sqrt{2}\sqrt{2\pi}}V_{1}\times\\ \int\limits_{0}^{t}d\tau{e^{-i\upsilon_{g}(k_{1}+k_{2})(t-\tau)}}{e^{-i\Omega_{1}\tau}}\left[{\beta(k_{2},\tau)+\beta(k_{1},\tau)}\right],\end{gathered}$ (9)
where we assume that at initial time $t=0$ no two-photon states are presented. Transforming this expression back to real-space using inverse of (7), we get:
$\begin{gathered}\gamma(x_{1},x_{2},t)=\\ -i\frac{V_{1}}{\sqrt{2}\upsilon_{g}}{e^{-i\Omega_{1}(t-\frac{x_{1}}{\upsilon_{g}})}}\beta\left({x_{2}-x_{1},t-\frac{x_{1}}{\upsilon_{g}}}\right)\theta(x_{1})\theta\left(t-\frac{x_{1}}{\upsilon_{g}}\right)\\ -i\frac{V_{1}}{\sqrt{2}\upsilon_{g}}{e^{-i\Omega_{1}(t-\frac{x_{2}}{\upsilon_{g}})}}\beta\left({x_{1}-x_{2},t-\frac{x_{2}}{\upsilon_{g}}}\right)\theta(x_{2})\theta\left({t-\frac{x_{2}}{\upsilon_{g}}}\right),\end{gathered}$ (10)
where we introduced a Heaviside step-function, $\theta(x>0)=1,\quad\theta(x<0)=0$. In order to avoid a jump at $x=0$, we apply standard regulation $\theta(x)=1/2$.
Next thing we need is to find $\beta(x,t)$, which then can be used to determine two-photon amplitude (10). Taking $\gamma(x_{1},x_{2},t)$ for $x_{2}=0$ and substitute it in (5), we obtain:
$\begin{gathered}\left({\frac{\partial}{\partial t}+{\upsilon_{g}}\frac{\partial}{{\partial x}}}\right)\beta(x,t)=-iV_{2}\delta(x){e^{i(\Omega_{1}-\Omega_{2})t}}\alpha(t)\\ -\frac{\Gamma_{1}}{2}\beta(x,t)-\Gamma_{1}F(x,t),\end{gathered}$ (11)
where we define decay rate of first-excited level as $\Gamma_{1}=V_{1}^{2}/\upsilon_{g}$, as well as some additional parameter:
$F(x,t)=e^{i\Omega_{1}\frac{x}{\upsilon_{g}}}\beta\left({-x,t-\frac{x}{\upsilon_{g}}}\right)\theta(x)\theta\left(t-\frac{x}{\upsilon_{g}}\right).$ (12)
To find explicit expression for $\beta(x,t)$, we use same procedure as for $\gamma(x_{1},x_{2},t)$. Multiplying (11) by $e^{-ikx}/\sqrt{2\pi}$ and integrating over $x$, we can use Fourier transform (7) to rewrite equation (11) in k-space. Then formal solution can be found as:
$\begin{gathered}\beta(k,t)=\beta_{0}(k){e^{-(ik\upsilon_{g}+\Gamma_{1}/2)t}}\\ -i\frac{V_{2}}{\sqrt{2\pi}}\int\limits_{0}^{t}d\tau{e^{-(ik\upsilon_{g}+\Gamma_{1}/2)(t-\tau)}}{e^{-i(\Omega_{2}-\Omega_{1})\tau}}\alpha(\tau)\\ -\Gamma_{1}\int\limits_{0}^{t}d\tau{e^{-(ik\upsilon_{g}+\Gamma_{1}/2)(t-\tau)}}F(k,\tau),\end{gathered}$ (13)
where $\beta_{0}(k)=\beta(k,0)$ is an amplitude for incident photon with frequency $\omega=k\upsilon_{g}$ at initial time $t=0$ and $F(k,t)$ is a Fourier transform of coefficient (12). Translating (13) back to real-space, we get:
$\begin{gathered}\beta(x,t)=e^{-\frac{\Gamma_{1}}{2}t}\beta_{0}(x-\upsilon_{g}t)-i\frac{V_{2}}{\upsilon_{g}}e^{-i\Delta_{\Omega}(t-\frac{x}{\upsilon_{g}})}e^{-\frac{\Gamma_{1}x}{2\upsilon_{g}}}\theta(x)\theta\left({t-\frac{x}{\upsilon_{g}}}\right)\alpha\left(t-\frac{x}{\upsilon_{g}}\right)\\ -\Gamma_{1}\theta(x)\theta\left(t-\frac{x}{\upsilon_{g}}\right)\int\limits_{0}^{x/\upsilon_{g}}dt^{\prime}{e^{-\frac{\Gamma_{1}}{2}t^{\prime}}}{e^{-i\Omega_{1}\left({t^{\prime}-\frac{x}{\upsilon_{g}}}\right)}}\beta\left({{\upsilon_{g}}t^{\prime}-x,t-\frac{x}{\upsilon_{g}}}\right),\end{gathered}$ (14)
where we introduce frequency between levels $\left|e\right\rangle$ and $\left|f\right\rangle$ as $\Delta_{\Omega}=\Omega_{2}-\Omega_{1}$ (see Fig. 1b). Change of upper limit in last integral is due to step function $\theta\left(x-\upsilon_{g}t^{\prime}\right)$.
Expression (14), aside from amplitude $\alpha(t-x/\upsilon_{g})$, is also contain shifted value $\beta(\upsilon_{g}t^{\prime}-x,t-x/\upsilon_{g})$, so it can’t be considered explicit. However, by performing a corresponding shift in (14), one can show that recursive part will cut off due to Heaviside functions, and remaining part will contain only terms proportional to incident single-photon field $\beta_{0}$ and amplitude of second-excited state $\alpha(t)$. Detailed calculations can be found in Appendix A. Thus, we obtain the explicit expression for a single-photon amplitude:
$\begin{gathered}\beta(x,t)=e^{-\frac{\Gamma_{1}}{2}t}\beta_{0}(x-\upsilon_{g}t)-i\frac{V_{2}}{\upsilon_{g}}e^{-i{\Delta_{\Omega}}(t-\frac{x}{\upsilon_{g}})}e^{-\frac{{{\Gamma_{1}}x}}{2\upsilon_{g}}}\theta(x)\theta\left({t-\frac{x}{\upsilon_{g}}}\right)\alpha\left({t-\frac{x}{\upsilon_{g}}}\right)\\ -\Gamma_{1}\theta(x)\theta\left({t-\frac{x}{\upsilon_{g}}}\right){e^{-\frac{\Gamma_{1}}{2}\left({t-\frac{x}{\upsilon_{g}}}\right)}}\int\limits_{0}^{x/{\upsilon_{g}}}dt^{\prime}{e^{-\frac{\Gamma_{1}}{2}t^{\prime}}}{e^{-i\Omega_{1}\left({t^{\prime}-\frac{x}{\upsilon_{g}}}\right)}}\beta_{0}\left[{{\upsilon_{g}}(t^{\prime}-t)}\right].\end{gathered}$ (15)
By taking (15) at point $x=0$, we can easily find the remaining amplitude of a second-excited state from (4):
$\begin{gathered}\alpha(t)=\alpha(0)e^{-\frac{\Gamma_{2}}{2}t}\\ -iV_{2}e^{-\frac{\Gamma_{2}}{2}t}\int\limits_{0}^{t}d\tau e^{\left({i\Delta_{\Omega}-\frac{\Gamma_{1}}{2}+\frac{\Gamma_{2}}{2}}\right)\tau}\beta_{0}(-\upsilon_{g}\tau),\end{gathered}$ (16)
where $\alpha(0)$ corresponds to initial state with fully excited transmon, and $\Gamma_{2}=V_{2}^{2}/\upsilon_{g}$ is a decay rate of second-excited state, similar to $\Gamma_{1}$.
Obtained expressions (10), (15) and (16) are the key results of this section. Taking particular initial state, we can derive explicit analytic expression for full wave function (3). These equations contain two possible initial conditions: 1) Transmon fully excited in state $\left|f\right\rangle$ and there are no free-photons in a waveguide; 2) Transmon is excited to state $\left|e\right\rangle$ and there is a single incident photon with arbitrary line shape $\beta_{0}\equiv\beta(k,0)$. First condition describes spontaneous emission of fully excited three-level system, and we analyze it below in details. Second one describes process of stimulated emission with added third level, and we left this task for future studies. It is also possible to solve this problem for initial state with two incident photons that scatters on a 3LE, but we omitted it to simplify the calculations.
IV Spontaneous decay of fully-excited transmon
IV.1 Solution in real-space
Let’s assume that initially there are no photons in a waveguide, and a three-level atom is in upper excited state $\left|f\right\rangle$:
$\alpha(0)=1,\qquad\beta_{0}(x)=0.$ (17)
In this case we have the process of spontaneous decay, which due to three levels results in emission of two photons.
Using initial condition (17) in (16), (15) and (10), we find exact expressions for all three probability amplitudes of a wavefunction:
$\alpha(t)=e^{-\frac{\Gamma_{2}}{2}t},$ (18a)
$\begin{gathered}\beta(x,t)=-i\sqrt{\frac{\Gamma_{2}}{\upsilon_{g}}}e^{-\left(i{\Delta_{\Omega}}+\frac{\Gamma_{2}}{2}\right)(t-x/\upsilon_{g})}e^{-\frac{\Gamma_{1}}{2}\frac{x}{\upsilon_{g}}}\times\\ \theta(x)\theta\left({t-\frac{x}{\upsilon_{g}}}\right),\end{gathered}$ (18d)
$\begin{gathered}\gamma(x_{1},x_{2},t)=-\sqrt{\frac{\Gamma_{1}\Gamma_{2}}{2\upsilon_{g}^{2}}}\times\\ \left({e^{-i\Omega_{1}\left(t-\frac{x_{1}}{\upsilon_{g}}\right)}e^{-\left(i\Delta_{\Omega}+\frac{\Gamma_{2}}{2}\right)\left(t-\frac{x_{2}}{\upsilon_{g}}\right)}e^{-\frac{\Gamma_{1}}{2}\frac{x_{2}-x_{1}}{\upsilon_{g}}}\Theta(x_{1},x_{2},t)}\right.\\ \left.{+{e^{-i{\Omega_{1}}\left({t-\frac{x_{2}}{\upsilon_{g}}}\right)}}e^{-\left({i\Delta_{\Omega}+\frac{\Gamma_{2}}{2}}\right)\left({t-\frac{x_{1}}{\upsilon_{g}}}\right)}e^{-\frac{\Gamma_{1}}{2}\frac{x_{1}-x_{2}}{\upsilon_{g}}}\Theta(x_{2},x_{1},t)}\right),\end{gathered}$ (18h)
where composite theta-function:
$\Theta(x_{1},x_{2},t)=\theta(x_{1})\theta(x_{2}-x_{1})\theta\left({t-\frac{x_{1}}{\upsilon_{g}}}\right)\theta\left({t-\frac{x_{2}}{\upsilon_{g}}}\right).$ (19)
Since here we work in a chiral approximation, all fields moves only right, so the emission of spontaneous photons only happens in forward direction, $x_{1,2}\equiv x>0$.
IV.2 State detection probabilities
Probability amplitudes (18), beside $\alpha(t)$, contain photonic dependence, proportional to real-space coordinate $x$ where photon is presumably located. So we can treat square root of those amplitudes as density of probability to find a photon in certain waveguide region. To find full probabilities to locate desired system state, one needs to integrate it over entire waveguide. By doing so, we obtain:
$P_{f,0}(t)=\left|\alpha(t)\right|^{2}=e^{-\Gamma_{2}t},$ (20a)
$P_{e,1}(t)=\int{dx}{\left|{\beta(x,t)}\right|^{2}}=\frac{\Gamma_{2}}{\Gamma_{2}-\Gamma_{1}}\left({e^{-\Gamma_{1}t}-e^{-\Gamma_{2}t}}\right),$ (20b)
$\begin{gathered}P_{g,2}(t)=\iint dx_{1}dx_{2}\left|\gamma(x_{1},x_{2},t)\right|^{2}=\\ =\frac{\Gamma_{1}\Gamma_{2}}{\Gamma_{2}-\Gamma_{1}}\left({\frac{1-e^{-\Gamma_{1}t}}{\Gamma_{1}}-\frac{1-e^{-{\Gamma_{2}}t}}{\Gamma_{2}}}\right).\end{gathered}$ (20e)
Figure 2: Probabilities to detect a certain state in entire waveguide, plotted with expressions (20): red solid line - transmon fully excited in level $\left|f\right\rangle$ and no photons presented; blue dashed line - transmon in first-excited state $\left|e\right\rangle$ and single photon in a waveguide; green dash-dotted line - transmon in a ground state $\left|g\right\rangle$, and two photons in a waveguide. Sum of all probabilities is equals to one and shown by thin dotted line. Different ratios of $\Gamma_{2}/\Gamma_{1}$ are shown above plots. Time is measured in $\Gamma_{2}$ units. Decay of upper level is constant, $\Gamma_{2}/\Omega_{1}=0.02$.
As one can see from (20), probabilities to detect a certain state of wavefunction (3) are independent on atomic resonant frequencies, and contain only decay rates $\Gamma_{1}$ and $\Gamma_{2}$. Plots of state-detection probabilities for a few $\Gamma_{2}/\Gamma_{1}$ ratios are presented in Fig. 2.
The decay of upper level $\left|f\right\rangle$ is not dependent on $\Gamma_{1}$, i.e. it’s stays the same for different $\Gamma_{2}/\Gamma_{1}$ ratios since we choose to take $\Gamma_{2}/\Omega_{1}=0.02$ constant (red solid line if Fig. 2). However, it is not true for the decay with an emission of single or two photons in the entire waveguide. There are two extreme cases: the decay of upper level is much larger, that of the middle one, $\Gamma_{2}/\Gamma_{1}\gg 1$, and inverse picture, when $\Gamma_{2}/\Gamma_{1}\ll 1$. For first case, the decay of upper levels is so fast, that transmon after short period of time starts to behave somewhat similar to a regular two-level system, formed by states $\left|e\right\rangle$ and $\left|g\right\rangle$ (see blue and green lines in Fig. 2a). For $\Gamma_{2}/\Gamma_{1}\leftarrow\infty$ first-excited state $\left|e\right\rangle$ become metastable. In second case decay of upper level $\left|f\right\rangle$ is almost instantly followed by decay of middle level $\left|e\right\rangle$, so, despite the cascaded process of emission, the two photons are produced nearly simultaneously (blue line in Fig. 2d is very close to zero, and green line is close to unity).
Typical transmon circuits have coupling ratio between first-excited and second-excited states as $g_{2}(k)/g_{1}(k)\sim\sqrt{3/2}$, see expressions (3.3-3.4) in Koch07 . The corresponding ratio of decay rates thus is $\Gamma_{2}/\Gamma_{1}\sim 1.5$, although some works mention very similar value of $\Gamma_{2}/\Gamma_{1}\sim 2$ Gasparinetti17 . We plot probabilities for common relation of $\Gamma_{2}/\Gamma_{1}=1.5$ in Fig. 2b. Albeit, recent progress in superconducting quantum circuits made possible creation of the so-called giant-atom transmon with highly tunable ratio $\beta=\Gamma_{2}/\Gamma_{1}\approx 0.3-60$ Vadiraj21 . However, further down we stick to more common transmon types with ratios of $\Gamma_{2}/\Gamma_{1}=1.5$ and $\Gamma_{2}/\Gamma_{1}=3$.
IV.3 Two-photon spectrum
By using Fourier transform (7), we can translate the real-space photon amplitudes (10) and (15) into a k-space of frequency modes:
$\beta(k,t)=i\sqrt{\frac{\Gamma_{2}\upsilon_{g}}{2\pi}}\frac{e^{-\left({i\omega+\frac{\Gamma_{1}}{2}}\right)t}-e^{-\left(i\Delta_{\Omega}+\frac{\Gamma_{2}}{2}\right)t}}{i(\omega-\Delta_{\Omega})-\frac{\Gamma_{2}-\Gamma_{1}}{2}},$ (21)
$\begin{gathered}\gamma(k_{1},k_{2},t)=\frac{\upsilon_{g}}{2\pi}\sqrt{\frac{\Gamma_{1}\Gamma_{2}}{2}}\frac{e^{-i(\omega_{1}+\omega_{2})t}}{i(\omega_{1}-\Delta_{\Omega})-\frac{\Gamma_{2}-\Gamma_{1}}{2}}\times\\ \left[{\frac{e^{\left({i\omega_{2}-i\Omega_{1}-\frac{\Gamma_{1}}{2}}\right)t}-1}{i\omega_{2}-i\Omega_{1}-\frac{\Gamma_{1}}{2}}-\frac{{{e^{\left({i\omega_{1}+i\omega_{2}-i\Omega_{2}-\frac{\Gamma_{2}}{2}}\right)t}}-1}}{i\omega_{1}+i\omega_{2}-i\Omega_{2}-\frac{\Gamma_{2}}{2}}}\right]\\ +\frac{\upsilon_{g}}{2\pi}\sqrt{\frac{\Gamma_{1}\Gamma_{2}}{2}}\frac{e^{-i(\omega_{1}+\omega_{2})t}}{i(\omega_{2}-\Delta_{\Omega})-\frac{\Gamma_{2}-\Gamma_{1}}{2}}\times\\ \left[{\frac{{e^{\left({i\omega_{1}-i\Omega_{1}-\frac{\Gamma_{1}}{2}}\right)t}}-1}{i\omega_{1}-i\Omega_{1}-\frac{\Gamma_{1}}{2}}-\frac{{e^{\left({i\omega_{1}+i\omega_{2}-i\Omega_{2}-\frac{\Gamma_{2}}{2}}\right)t}}-1}{i\omega_{1}+i\omega_{2}-i\Omega_{2}-\frac{\Gamma_{2}}{2}}}\right],\end{gathered}$ (22)
where in $\beta(k,t)$ photon frequency $\omega$ corresponds to first emitted photon.
Note that if we take $t\rightarrow\infty$, single photon amplitude (21), as well as $\alpha(t)$, will be equals to zero. However, two-photon amplitude (22) for large times takes non-zero value. It is logical prediction, because final state of excited three-level system is obviously a two photons in a waveguide with the transmon being in the ground state.
Let’s define the spectral density of spontaneous emission of initially excited transmon as follows:
$\begin{gathered}S_{sp}(\omega_{1},\omega_{2})=\upsilon_{g}^{2}{\left|{\gamma(k_{1},k_{2},t\to\infty)}\right|^{2}}=\\ =\frac{1}{8\pi^{2}}\frac{\Gamma_{1}\Gamma_{2}}{{{\left(\omega_{1}+\omega_{2}-\Omega_{2}\right)}^{2}+\frac{\Gamma_{2}^{2}}{4}}}\times\\ \frac{{{\left(\omega_{1}+\omega_{2}-2\Omega_{1}\right)}^{2}}+\Gamma_{1}^{2}}{\left[{{{\left({\omega_{1}-\Omega_{1}}\right)}^{2}}+\frac{\Gamma_{1}^{2}}{4}}\right]\left[{{{\left({\omega_{2}-\Omega_{1}}\right)}^{2}}+\frac{\Gamma_{1}^{2}}{4}}\right]}.\end{gathered}$ (23)
We plot this two-photon density in Fig. 3.
Unlike the probabilities (20), spectral density (23) depends not only on decay rates $\Gamma_{1},\Gamma_{2}$, but also on resonant frequencies of 3LE. It is typical to describe energy-level spacing of a transmon via an anharmonicity parameter (Koch07, ):
$\alpha_{r}=\frac{\Delta_{\Omega}-\Omega_{1}}{\Omega_{1}}=\frac{\Omega_{2}-2\Omega_{1}}{\Omega_{1}}\approx-{\left(\frac{8E_{J}}{E_{C}}\right)^{-1/2}},$ (24)
where $E_{J}$ is a Josephson energy, and $E_{C}$ is a charging energy - common characteristics of superconducting quantum circuits Blais21 . Subscript $r$ stands for ’relative’, thus expression (24) is dimensionless and shows relation of energies between states $\left|f\right\rangle,\;\left|e\right\rangle$ and energies between levels $\left|e\right\rangle,\;\left|g\right\rangle$. In literature it is also common to use a non-relative anharmonicity $\alpha=(\Delta_{\Omega}-\Omega_{1})=\alpha_{r}\Omega_{1}$, that is measured in hertz, but here we choose to use only relative parameter to avoid confusion with probability amplitude $\alpha(t)$. Typical experimental values of $\alpha$ usually falls within the range 100-300 MHz Krantz19 , or in relative units from $2\%$ up to $6\%$. Note that anharmonicity for superconducting qubits are always negative, due to smaller spacing between upper levels.
Figure 3: Two-photon spectral density of spontaneous emission from three-level transmon $S_{sp}(\omega_{1},\omega_{2})\Gamma_{2}^{2}$, plotted using expression (23). (a) high anharmonicity $\alpha_{r}=-5\%$ and typical decay rates relation $\Gamma_{2}/\Gamma_{1}=3/2$; (b) smaller anharmonicity $\alpha_{r}=-3\%$, and $\Gamma_{2}/\Gamma_{1}=3/2$; (c) faster decay of upper level $\Gamma_{2}/\Gamma_{1}=3$ and $\alpha_{r}=-3\%$; (d) same decay $\Gamma_{2}/\Gamma_{1}=3$ and smaller anharmonicity $\alpha_{r}=-2\%$. Decay of upper level taken constant $\Gamma_{2}/\Omega_{1}=0.02$, except for (e), where weak-coupling is analyzed for $\Gamma_{2}/\Omega_{1}=0.001$, $\Gamma_{2}/\Gamma_{1}=1.5$ and $\alpha_{r}=-3\%$.
Figure 4: Spectral density of two identical photons (25) vs frequency detuning - (a) and (b), and vs detuning and anharmonicity - (c) and (d). (a) and (c) are plotted for $\Gamma_{2}/\Gamma_{1}=3/2$ ratio, and (b) and (d) for $\Gamma_{2}/\Gamma_{1}=3$. In (a), (b) solid magenta line corresponds to $\alpha_{r}=-3\%$, green dashed line to $\alpha_{r}=-4\%$, and black dotted line to $\alpha_{r}=-6\%$. Vertical grey lines show points for $\delta_{\omega}/\Omega_{1}=\alpha_{r}/2$. In (c), (d) thin white lines indicate resonant frequency $\omega=\Omega_{1}$. Decay of upper level taken constant $\Gamma_{2}/\Omega_{1}=0.02$.
From Fig. 3 we see that spectral density of spontaneous emission have two distinct peaks, which are a result of non-distinguishability of photons. Each peak corresponds to one resonant photon with frequency $\omega=\Omega_{1}$ and second photon with smaller frequency $\omega=(1+\alpha_{r})\Omega_{1}$ (remember that $\alpha_{r}$ is negative). But because both photons are indistinguishable, we can’t exactly say which photon is ’first’, and which is ’second’, thus two peaks appear. For bigger anharmonicity, two peaks are further apart, because of change of transition frequencies in 3LE. If we implement weak-coupling, $\Gamma/\Omega_{1}=0.001$ (in practical circuits its usually $\Gamma<5$ MHz), both peaks are tighten into almost point-like shapes (see Fig. 3e). That means that we have two photons with distinct frequencies, correspondent to atom-transitions, $\Omega_{1}$ and $(1+\alpha_{r})\Omega_{1}$. This behavior is in good agreement with experimental work by Gasparinetti et al Gasparinetti17 .
However, if we take stronger coupling, typical for circuit QED systems (for example, $\Gamma_{2}\sim 100$ MHz), we see the appearance of frequency correlation between the two photons. This is expressed by the appearance of third peak at resonant frequency $\omega/\Omega_{1}=1$, which is start to become visible for $\Gamma_{2}/\Gamma_{1}=1.5$ in Fig. 3(b), and more pronounced for $\Gamma_{2}/\Gamma_{1}=3$ in Fig. 3(c)-(d). Thus, even if three-level system have non-symmetrical level structure, there is a probability that during spontaneous decay emitted photons would have same frequency $\omega_{1}=\omega_{2}=\Omega_{1}$. This effect is connected to interference between first emitted photon and 3LE being still excited in state $\left|e\right\rangle$, which is arise due to almost equidistant level structure of low-anharmonicity transmon and rather strong coupling between the artificial atom and the field modes.
IV.4 Emission spectra for identical photons
Let’s study spectral density (23) for identical photons in closer detail. Taking $\omega_{1}=\omega_{2}$ we get:
$S_{sp}(\omega_{1}=\omega_{2})=\frac{1}{2\pi^{2}}\frac{\Gamma_{1}}{\delta_{\omega}^{2}+\frac{\Gamma_{1}^{2}}{4}}\times\frac{\Gamma_{2}}{(2\delta_{\omega}-\alpha_{r}\Omega_{1})^{2}+\frac{\Gamma_{2}^{2}}{4}},$ (25)
where we introduce detuning from resonant frequency for both photons as $\delta_{\omega}=\omega-\Omega_{1}$.
Function (25) have two maximums: first one corresponds to frequency between lower levels $\Omega_{1}$ and thus $\delta_{\omega}=0$, and second one corresponds to decay of upper level with anharmonicity shift, $\delta_{\omega}=-\alpha_{r}\Omega_{1}/2$. Both these peaks can be observed in Fig. 4(a)-(b), for decay rates $\Gamma_{2}/\Gamma_{1}=3/2$ and $\Gamma_{2}/\Gamma_{1}=3$. Two-dimensional plot of (25), showing how spectral density also changes with anharmonicity, presented in Fig. 4(c)-(d).
The maximum of spectral density for identical photons should be for $\alpha_{r}\rightarrow 0$, i.e. in the case of equidistant levels with $\Omega_{2}=2\Omega_{1}$. However, as can be seen from Fig. 4, frequencies of two photons not always have to be resonant with $\Omega_{1}$. For example, if ratio between decay rates is $\Gamma_{2}/\Gamma_{1}=1.5$, like in Fig. 4(a), emitted photons both would have maximum spectral density on frequency, close to $\Omega_{1}(1+\alpha_{r}/2)$. This also can be observed on two-frequency graph Fig. 3(b) in diagonal cross-section. Thus, by changing both decay rates $\Gamma_{1},\Gamma_{2}$ and anharmonicity parameter $\alpha_{r}$ of excited three-level transmon, it is possible to create correlated photon pairs with same frequencies. Superconducting circuits are very well suitable for this task, since they allow tuning of couplings and anharmonicity parameters either by sample production techniques or via application of external magnetic fields.
V Conclusion
In summary, we present a theoretical description of spontaneous decay of a fully-excited three-level system in a one-dimensional open waveguide. Using the real-space description of the system, we obtain non-stationary analytical expressions for the probability amplitudes for the whole system. We show that the dynamics of spontaneous emission is heavily dependent on the ratio $\Gamma_{2}/\Gamma_{1}$, where $\Gamma_{2}$ ($\Gamma_{1}$) is the radiative decay rate of the second (first) excited state. Transferring the obtained amplitudes into the momentum space allowed us to study the spectral density of emitted photons, from which we discover the appearance of frequency correlation for strong coupling. It is expressed in the creation of two photons with identical frequencies, even if the three-level atom structure is anharmonic. The frequency correlation is stronger for smaller anharmonicity, i.e., when the levels are close to equidistant.
More detailed analysis showed that the spectral density of identical photons can have two maxima. The first one corresponds to the resonant frequency of the lower levels $\omega=\Omega_{1}$ and prevails when the decay rates ratio is rather high, $\Gamma_{2}/\Gamma_{1}\approx 3$. The second maximum, however, depends on the anharmonicity parameter as $\omega\approx\Omega_{1}(1+\alpha_{r}/2)$, and it dominates for $\Gamma_{2}/\Gamma_{1}\approx 1.5$, which is a rather common value for typical transmon circuits. Thus, by adjusting these parameters one can obtain pairs of identical photons with tunable frequency, which can be utilized for quantum information and communication applications.
Throughout our analysis we take specific parameters that are common for superconducting transmon qubits, coupled to a microwave coplanar waveguide. However, our results are obtained in a general form, and can be used to describe other physical systems of three-level emitters in 1D waveguides, like Rydberg atoms or quantum dots in nanoplasmonic waveguides.
Here we describe only the process of spontaneous emission of a fully-excited three-level emitter. However, the obtained solutions in real-space (10), (15) and (16) allow one to study also the case of an incident photon pulse, scattering on a half-excited three-level system in state $\left|e\right\rangle$. Moreover, our expression (15) is valid for an arbitrary pulse shape $\beta_{0}(x)$, which can be, for example, gaussian or rectangular. This task is quite interesting from the point of view of stimulated emission, and can be addressed in future studies.
Acknowledgements.
The authors thank A. N. Sultanov for fruitful discussions and helpful comments on the manuscript. The work is supported by the Ministry of Science and Higher Education of Russian Federation under the project FSUN-2026-0004. O. Chuikin and O. Kibis acknowledge the financial support from the Foundation for the Advancement of Theoretical Physics and Mathematics “BASIS”.
Appendix A Derivation of single-photon amplitude (15)
We start by incorporating the following shift:
$x\to{\upsilon_{g}}t^{\prime}-x;\qquad t\to t-\frac{x}{\upsilon_{g}};$ (26)
and using it in (14). By doing so we obtain:
$\beta\left({{\upsilon_{g}}t^{\prime}-x,t-\frac{x}{{{\upsilon_{g}}}}}\right)={A_{0}}-\theta({\upsilon_{g}}t^{\prime}-x)\theta(t-t^{\prime})\left[{{A_{1}}+{A_{2}}}\right],$ (27)
where for simplicity we introduce some coefficients as follows:
$A_{0}$ $=e^{-\frac{\Gamma_{1}}{2}(t-x/\upsilon_{g})}\beta_{0}\left[{{\upsilon_{g}}(t^{\prime}-t)}\right],$ (28a)
$A_{1}$ $={\Gamma_{1}}\int\limits_{0}^{t^{\prime}-x/{\upsilon_{g}}}dt^{\prime\prime}{e^{-\frac{\Gamma_{1}}{2}t^{\prime\prime}}}{e^{-i\Omega_{1}\left({t^{\prime\prime}-t^{\prime}+\frac{x}{\upsilon_{g}}}\right)}}\beta\left({x,t-t^{\prime}}\right),$ (28b)
$A_{2}$ $=i\frac{V_{2}}{\upsilon_{g}}{e^{-i{\Delta_{\Omega}}(t-x/{\upsilon_{g}})}}{e^{(i{\Delta_{\Omega}}-{\Gamma_{1}}/2)(t^{\prime}-x/{\upsilon_{g}})}}\alpha\left({t-t^{\prime}}\right).$ (28c)
Substituting (27) back to (14), we get:
$\begin{gathered}\beta(x,t)={e^{-\frac{\Gamma_{1}}{2}t}}{\beta_{0}}(x-{\upsilon_{g}}t)-i\frac{V_{2}}{\upsilon_{g}}{e^{-i{\Delta_{\Omega}}(t-x/{\upsilon_{g}})}}{e^{-\frac{{\Gamma_{1}}x}{2\upsilon_{g}}}}\theta(x)\theta\left({t-\frac{x}{\upsilon_{g}}}\right)\alpha\left({t-\frac{x}{\upsilon_{g}}}\right)\\ -{\Gamma_{1}}\theta(x)\theta\left({t-\frac{x}{\upsilon_{g}}}\right)\int\limits_{0}^{x/{\upsilon_{g}}}{dt^{\prime}}{e^{-\frac{\Gamma_{1}}{2}t^{\prime}}}{e^{-i{\Omega_{1}}\left({t^{\prime}-\frac{x}{\upsilon_{g}}}\right)}}\left({{A_{0}}-\theta({\upsilon_{g}}t^{\prime}-x)\theta(t-t^{\prime})\left[{{A_{1}}+{A_{2}}}\right]}\right).\end{gathered}$ (29)
The terms, proportional to $\int_{0}^{x/\upsilon_{g}}{\theta({\upsilon_{g}}t^{\prime}-x)dt^{\prime}}$ will drop out due to mismatch of intervals, since the condition of the theta function is $t’>x/\upsilon_{g}$. Thus, only term containing $A^{(0)}$ remain, proportional to the initial field $\beta_{0}$. Rewriting (29) in explicit form, we get the final expression (15).